OpenStudy (anonymous):

Sandi can make a quilt in 9 days and eva can make a quilt in 6 days. In how many days can they make a quilt working together?

OpenStudy (wolf1728):

There is a formula for this: (S x E) / (s+e) 9 x 6 / 9 +6 = 54 / 15 Equals 3.6 days

OpenStudy (anonymous):

thank you so much!!! i appreciate your help!

OpenStudy (wolf1728):

and thank you

OpenStudy (anonymous):

1 / (1/6 + 1/9) = 3.6 days

OpenStudy (wolf1728):

Okay to see it worked out Sandy can make 1/9 of a quilt in one day Eva can make 1/6 of a quilt in one day. Both working together they make (1/6 + 1/9) part of a quilt in one day. (1/6 + 1/9) = (3/18 + 2/18) = 5/18 of a quilt per day. So, a whole quilt takes 1/(5/18) days which equals (18/5) which equals 3 + 3/5 days (or 3 + 6/10 days)

OpenStudy (wolf1728):

Gee - 3 medals for one question. Thanks folks.